what temp in celsius is 1.5 times that of Fahrenheit?
Answer by Alan3354(69443)

C = 1.5F = 1.5*(C*1.8 + 32)
C = 2.7C + 48
C = -48/1.7 =~ -28.235 degs
-----------------